Small screws are stored 50 per box, in boxes measuring 3 x shelves. The shelf dimensions are 1 (high), 1 (deep), and shelv
Answer #1

a) No. of boxes of screws

800,000 / 50 = 16,000 Boxes

Each shelf is 12" x 12" x 36" (hxdxw) and each box is 3" x 3" x 3"

hence each shelf can take

(12 x 12 x 36) / (3 x 3 x 3) = 192 boxes

i.e 4 layers stacked, with each layer consisting of 4 rows of 12 boxes each

Therefore the number of shelves needed is

16,000/192 83.33 ︵ะ 84

hence a minimum of 84 shelves are required

b) Each units consists of 10 shelves; so we need

84/10 = 8.4 ะ 9

9 units are required

c) as each unit is 3 ft wide, we need a minimum of

3 x 9 = 27 aisle feet